# Edge of Stochastic Stability: Revisiting the Edge of Stability for SGD

### v3

@@ 3 SGD Typically Occurs at the EoSS
Extending Cohen et al. (2021), we identify a stable quantity in mini-batch SGD: Batch Sharpness. We empirically demonstrate that SGD operates in a regime we term Edge of Stochastic Stability, observed across multiple datasets and architectures. Batch Sharpness emerges as the key quantity governing this regime, serving as the analogous for mini-batch training of the sharpness.
@@
Figure 3: Comparing different sharpness measures. Red: observed sharpness on the step’s mini batch on the batch used for the step; Green: average Batch Sharpness (Definition 1); Blue: full-batch $\lambda_{\max}$. Top row: MLP (2 hidden layers of width 512); middle: 5-layer CNN; bottom: ResNet-14; all trained on an 8k subset of CIFAR-10.
3.1 Mini-Batch Sharpness
Let $L_{B}(\theta)\!=\!\frac{1}{b}\!\sum_{(x_{i},y_{i})\in B}\!\tilde{L}\bigl{(}f_{\theta}(x_{i}),y_{i}\bigr{)}$ be the empirical loss on a randomly sampled batch $B$ of size $b$. Denote by $\mathcal{H}(L_{B})$ the Hessian w.r.t. $\theta$ of $L_{B}$, and let $\nabla L_{B}(\theta)$ be its gradient. We define:
Definition 1 (Batch Sharpness).
Let $\mathcal{P}_{b}$ be the distribution of random batches of size $b$. The Batch Sharpness is
Batch $\text{Batch Sharpness}\;=\;\frac{\mathbb{E}_{B\sim\mathcal{P}_{b}}\bigl{[}\nabla L_{B}^{\top}\,\mathcal{H}(L_{B})\,\nabla L_{B}\bigr{]}}{\mathbb{E}_{B\sim\mathcal{P}_{b}}\bigl{[}\|\nabla L_{B}\|^{2}\bigr{]}}\,.$
Denote by $\lambda_{\max}:=\lambda_{\max}\left[\mathcal{H}(L)\right]$ the largest eigenvalue of the Hessian of the full-batch loss. While Batch Sharpness behaves as $\lambda_{\max}$ in EoS, the two quantities differ fundamentally. We explore this distinction further in Section 4.
3.2 The Edge of Stochastic Stability
1. Stabilization of Batch Sharpness:
We establish that SGD operates in a stability regime analogous to EoS, which we term Edge of Stochastic Stability. In this regime Batch Sharpness stabilizes at $2/\eta$: {mdframed}
SGD tends to train in an instability regime we call Edge of Stochastic Stability. Precisely, after a phase of progressive sharpening, Batch Sharpness reaches a stability level of $2/\eta$, and hovers there.
2. Stabilization of $\lambda_{\max}$:
Also $\lambda_{\max}$ typically stabilizes. However, it does so at lower values that depends on the trajectory taken (Figures 3 and 5). We discuss the stabilization level of $\lambda_{\max}$ in Section 3.3 and 5. Consequently, standard EoS analysis that relies solely on $\lambda_{\max}$ cannot explain SGD’s actual training behavior in the small-batch setting.
3. Catapults:
Unlike EoS, EoSS stabilizes the expectation of a quantity which the algorithm sees one observation at time. Occasionally, a series of batches sampled have outlying sharpness–that is too high for the stable regime–and steps overshoot, triggering a catapult effect, where Batch Sharpness spikes before rapidly decreasing (Figure 4). This is typically followed by renewed progressive sharpening, returning to the EoSS regime. This results in a catapult phase for the training loss, aligning with, and maybe explaining, previous observations about catapult behaviors, see, e.g., Lewkowycz et al. (2020); Zhu et al. (2024).
Figure 4: Catapults during EoSS. During EoSS, Batch Sharpness may through cycles of progressive sharpening and stabilizations. Red dots are the observed sharpnesses of the batches used for the step, the green line is Batch Sharpness.
@@ 3. Catapults:
Figure 5: Effects of changing step size or batch size in EoSS. (a) Increasing the learning rate causes a catapult spike before Batch Sharpness settles at the new $2/\eta$. (b) Decreasing the learning rate prompts renewed progressive sharpening. (c) Increasing the batch size narrows the gap between Batch Sharpness and $\lambda_{\max}$. (d) Decreasing the batch size. Triangles mark the steps at which the hyperparameters were changed.
3.3 Batch Sharpness Governs EoSS
Following Cohen et al. (2021) we show here how the dynamics of Batch Sharpness changes when hyper-parameters shift along the dynamics. Interestingly, while $\lambda_{\max}$ and other full-batch Hessian metrics may stabilize, we show here their specific plateaus or trajectories depend on initial conditions, not only on the hyperparameter. By contrast, $\lambda_{\max}(\mathcal{H}(L_{B}))$—the top eigenvalue of the mini-batch Hessian—does consistently stabilizes and shift with hyperparameters.
Changing the step size.
Increasing the step size $\eta$ triggers a catapult spike in all the quantities in considerations and the training loss, before Batch Sharpness re-stabilizes near the updated threshold $2/\eta$ (Fig. 5(a)). This therefore pushes $\lambda_{\max}$ lower. Crucially, if stability was governed solely by $\lambda_{\max}$, this step-size adjustment would have left the dynamics unchanged, as $\lambda_{\max}$ remains below both the original and revised $2/\eta$ thresholds. Conversely, reducing $\eta$ (Fig. 5(b) raises the $2/\eta$ threshold and prompts a new phase of progressive sharpening of Batch Sharpness. Here, $\lambda_{\max}$ also rises, but ultimately stabilizes at a lower value than if the entire training had run with the smaller step size. Again, if stability was governed solely by $\lambda_{\max}$, this step-size adjustment would have had the same effect as starting from scratch with the new step size. Notably, Batch Sharpness initially decreases—a counterintuitive result potentially linked to shifting alignments between gradients and Hessian eigenvectors. This behavior warrants further investigation.
Changing the batch size.
Increasing the batch size reduces the gap between Batch Sharpness and $\lambda_{\max}$, therefore making Batch Sharpness lower,(Fig. 5(c)), and thus below the $2/\eta$ threshold. This causes a new progressive sharpening phase until Batch Sharpness again reaches the $2/\eta$ level, making $\lambda_{\max}$ stabilize at a new, higher level, again lower than the one of the training started from scratch with the new batch size. This shows that it is indeed the Batch Sharpness that is the "limiting" quantity in EoSS. Conversely, decreasing the batch size induces a brief catapult phase, after which Batch Sharpness returns to its previous plateau, as expected. This forces full-batch $\lambda_{\max}$ to adjust to a lower level.
Overall, we find that Batch Sharpness governs EoSS behavior—mirroring how $\lambda_{\max}$ operates in the full-batch EoS—while the actual full-batch $\lambda_{\max}$ lags behind or settles inconsistently, underlining the mini-batch nature of SGD stability.
4 Mathematics of Stability
It may seem surprising at first that Batch Sharpness is not a quantity of the same nature as $\lambda_{\max}$ or is not a function of the full-batch Hessian. The reason is deep into the nature of mini-batch SGD. After recalling the computations of Cohen et al. (2021) in Section 4.1, we show what is the setting of mini-batch SGD and what is the notion of stability which depends on Batch Sharpness in Section 4.2. Later in Section 4.3 we motivate why the stability of mini-batch SGD, surprisingly, can not depend on full-batch quantities.
4.1 A Quadratic Objective: Full-Batch GD
We recall here the arguments of Cohen et al. (2021) and Lee & Jang (2023) for quadratic objective functions to later show how stability differs between the two settings.
Assume we fix a learning rate $\eta$ for the whole training. Consider the example of the quadratic Taylor approximation of a loss function:
$f(x)=\frac{1}{2}x^{\top}\mathbf{A}x+\mathbf{b}^{\top}x+\mathbf{c},$
where $\mathbf{A}$ is the Hessian matrix (symmetric and positive semidefinite), $\mathbf{b}$ is the gradient vector, and $\mathbf{c}$ is a constant. Here, as Cohen et al. (2021) observed, when $\eta\lambda_{\max}(\mathbf{A})>2$ the component of $x_{t}-x^{*}$ parallel to the eigenvector relative to $\lambda_{\max}(\mathbf{A})$ explodes with a power law. This is equivalent to imposing the inequality (1) at every step.
[…]
Indeed, eventually, we have $f(x_{t+1})>f(x_{t})$, i.e., if and only if $\lambda_{\max}\leq 2/\eta$.
4.2 Quadratic Objectives: Mini-Batch GD
Empirically, progressive sharpening halts when the mini-batch analog of (1) reaches zero in expectation. Concretely, EoSS arises if, at each step,
$\eta\,\|\nabla f_{B}(x_{t})\|^{2}\;\sim\;\frac{\eta^{2}}{2}\,\nabla f_{B}(x_{t})^{\top}\mathbf{A}_{B}\nabla f_{B}(x_{t}),$ (2)
preventing the single-batch loss from diverging. This makes sense, as the gradient step is on the mini-batch loss, thus the training is stuck in an oscillatory regime when the single batch loss oscillates (on average) [^1] After completion of this manuscript we found Lee & Jang (2023). We compare this with their work in Appendix C..
One might expect this to imply $\lambda_{\max}(\mathcal{H}(L_{B}))\leq 2/\eta$, as in full-batch (1), but that is typically not the case. What is inducing (1) to rely on $\lambda_{\max}$ is that the Hessians at every step share the eigenvectors relative to the highest eigenvalue. In mini-batch SGD, the sampled Hessians $\mathbf{A}_{B}$ generally have completely different eigenspaces than $\mathbf{A}$ and between each others. Indeed, the case of mini-batch SGD, we are working with the following quadratic Taylor approximation:
$f_{B}(x)=\frac{1}{b}\sum_{i\in B}\left(\frac{1}{2}x^{\top}\mathbf{A}_{i}x+\mathbf{b}_{i}^{\top}x+\mathbf{c}_{i}\right),$ (3)
where $B\subset\{1,2,\ldots,n\}$ is a randomly sampled mini batch of size $b$, and $\mathbf{A}_{i}$, $\mathbf{b}_{i}$, and $\mathbf{c}_{i}$ represent coefficients on the quadratic losses of individual data points. Here the full-batch Hessian $\mathbf{A}$ is given by $\mathbf{A}=\frac{1}{n}\sum_{i=1}^{n}\mathbf{A}_{i}$. Even if $\lambda_{\max}(\mathbf{A})>2/\eta$ with eigenvector $v_{\max}$, certain $\mathbf{A}_{i}$ may map $v_{\max}$ to 0 or a completely different vector, avoiding the blow-up seen in the full-batch regime.
Enforcing (2) in this setting (with $\mathbf{b}_{i}=0$) thus results in
$\frac{2}{\eta}\ \geq\ \sup_{v\in\mathbb{S}^{d}}\ \frac{v^{\top}\mathbb{E}_{i}\left[(\mathbf{A}_{i})^{3}\right]v}{v^{\top}\mathbb{E}_{i}\left[(\mathbf{A}_{i})^{2}\right]v}\ \begin{subarray}{c}\text{ which does not}\\ \text{simplify to}\end{subarray}\ \lambda_{\max}(\mathbf{A}).$ (4)
4.3 Why Stability cannot Depend on Full-Batch
Let us now look at possible full-batch stability results, as developed by Cohen et al. (2021). If the right stability notion for mini-batch SGD depended on the full-batch Hessian or gradients, then it would be independent on $\gamma$, and this can not be the case. For instance, if a notion as the one in Equation (1) would explain stability, then $\alpha$ could grow until $\alpha=2/\eta$. However, no matter the size of $\alpha$, if $\gamma>\max\big{\{}\alpha,\tfrac{2}{\eta}-\alpha\big{\}}$ a mini-batch iteration $x_{t+1}=(1-\eta\mathbf{A}_{i})x_{t}=$ increases the size of the second component of $x$ by a multiplicative factor bigger than 1.
Thus a study of the stability of the system can not depend on the full-batch Hessian but has to depend on how big the oscillations due to the size of $\gamma$ are, i.e., on the higher moments of the distribution of the mini-batch Hessian. Note that this situation would be even more extreme if we had (as in the practice of deep learning) the top eigenvectors of the mini-batch Hessians to point in completely different directions, not just to have high variance. As a sanity check, any stability threshold dependent on the higher moments of the mini-batch Hessian or gradients–as Batch Sharpness or $\mathbb{E}_{B}[\lambda_{\max}(\mathcal{H}(L_{B})]$–would induce $\eta_{\max}$ to depend on $\gamma$ too.
5 The Geometry of the Landscape at EoSS
We devote this section to the discussion of the sizes during training of different quantities which may be or interest of practitioners of theoreticians, as the $\lambda_{\max}$ of full-batch and mini-batch Hessians in terms of learning rate, batch size, and critical batch size. For a batch size $b$, denote by
$\lambda_{\max}^{b}:=\mathbb{E}_{B\sim\mathcal{P}_{b}}\big{[}\lambda_{\max}(\mathcal{H}(L_{B}))\big{]}.$

### v4

@@ 4 SGD Typically Occurs at the EoSS
We introduce in this section Batch Sharpness, our notion of implicit curvature. We proceed characterizing the EoSS phenomenon, and establishing it empirically. We empirically find that when Type-2 oscillations occur, Batch Sharpness aligns closely with $2/\eta$. Interestingly, throughout most of the training—even during progressive sharpening—also Type-1 oscillations are present.
4.1 Mini-Batch Sharpness
Throughout this paper $B\subset\mathcal{D}$ denotes a random mini-batch of size $b$ drawn from a fixed sampling distribution $\mathcal{P}_{b}$. For model parameters $\theta\in\mathbb{R}^{d}$ let $L_{B}(\theta)=\frac{1}{b}\sum_{(x_{i},y_{i})\in B}\!\tilde{L}\!\bigl{(}f_{\theta}(x_{i}),\,y_{i}\bigr{)}$, $L(\theta)=\mathbb{E}_{B\sim\mathcal{P}_{b}}\!\bigl{[}L_{B}(\theta)\bigr{]}$ be the mini-batch and full-batch losses, respectively. Write $\mathcal{H}(L_{B})=\nabla_{\theta}^{2}L_{B}(\theta)$.
Definition 1 (Batch Sharpness).
We define Batch Sharpness as the ratio
Batch $\textit{Batch Sharpness}\,(\theta)\quad:=\quad\frac{\mathbb{E}_{B\sim\mathcal{P}_{b}}\bigl{[}\nabla L_{B}(\theta)^{\top}\,\mathcal{H}(L_{\mathbf{B}})\,\nabla L_{B}(\theta)\bigr{]}}{\mathbb{E}_{B\sim\mathcal{P}_{b}}\bigl{[}\|\nabla L_{\mathbf{B}}(\theta)\|^{2}\bigr{]}}\,.$ (2)
Interpretation.
For each mini-batch, $\nabla L_{B}$ is the stochastic search direction produced by SGD. Thus the normalized version of $\nabla L_{B}^{\top}\,\mathcal{H}(L_{B})\,\nabla L_{B}$, Batch Sharpness, measures the expected directional curvature of the mini-batch loss surface along the corresponding mini-batch gradients. For locally quadratic losses a constant step $\eta$ is, on average, stable on the mini-batch landscape iff Batch Sharpness< $\textit{Batch Sharpness}\,<\,2/\eta$, see Lemma 1; crossing the threshold induces the catapult–type divergence that marks the EoSS regime, see Section 5.
Relation to earlier notions.
Lee and Jang [17] showed that the ratio $\nabla L^{\top}\mathcal{H}\nabla L/\|\nabla L\|^{2}$ settles at $2/\eta$ during EoS and coincides with $\lambda_{\max}(\mathcal{H})$ only when the gradient aligns with the top eigenvector. Batch Sharpness extends that directional viewpoint to the stochastic setting by replacing $(\nabla L,\mathcal{H})$ with their mini-batch counterparts $\bigl{(}\nabla L_{B},\mathcal{H}(L_{B})\bigr{)}$ before taking expectations [^6] Importantly, see Appendix L, the right notion of curvature/stability for mini-batch algorithms has to depend on different statistics or moments of the mini-batch Hessians, not simply on the average of the Hessians as in full-batch. Batch Sharpness, interestingly, translates to a ratio between third and second moment in the quadratic setting..
4.2 The Edge of Stochastic Stability
We characterize here the phenomenon of the Edge of Stochastic Stability. We verify the emergence of EoSS across of a range of step sizes, batch sizes and architectures (Figure 6 and Appendix M); datasets (CIFAR-10 and SVHN, Appendix N); and dataset sizes (8k and 32k subsets, Figure 7).
1. Stabilization of Batch Sharpness.
SGD typically operates in a regime analogous to EoS: {mdframed}
SGD tends to train in a regime we call Edge of Stochastic Stability. Precisely, after a phase of progressive sharpening, Batch Sharpness reaches a stability level of $2/\eta$, and hovers there.
In particular, the level of plateau of Batch Sharpness is $2/\eta$ independent of the batch size (Figure 1). Importantly, Type-1 oscillations happen throughout most of the training as highlighted by the quantity of Proposition 1, see Figure 4, but they do not impact progressive sharpening which leads to the second phase of EoSS stabilization and Type-2 oscillations. Importantly, analogously to EoS, training continuous and the loss continues to decrease while Batch Sharpness is constrained by the learning rate size.
2. Stabilization of $\lambda_{\max}$ and GNI.
Crucially, stabilization of Batch Sharpness around $2/\eta$ happens while GNI is stabilized at $2/\eta$ and induces a corresponding stabilization of $\lambda_{\max}$. However, $\lambda_{\max}$ consistently settles at a lower level, due to a batch-size–dependent gap between the two (Figures 6). This is also influenced by the specific optimization trajectory (Figure 7). We make observations on this relationship and the factors determining the stabilization gap in Section 6.
3. Catapults.
Figure 5: Catapults at EoSS. During EoSS, Batch Sharpness goes through cycles of progressive sharpening and stabilizations. Notations follow Figure 6.
Unlike in EoS, in the EoSS regime what is stabilized is the expectation of a quantity which the algorithm sees one observation at time. Occasionally, a sequence of sampled batches exhibits anomalously high sharpness–that is too high for the stable regime–and steps overshoot, triggering a catapult effect, where Batch Sharpness spikes before rapidly decreasing (Figure 5). This is typically followed by renewed progressive sharpening, eventually returning to the EoSS regime. This results in a catapult phase for the training loss, aligning with, and maybe explaining, previous observations about catapult behaviors, e.g., [16, 50].
@@ 4.2 The Edge of Stochastic Stability
Figure 6: Comparing different sharpness measures. Red: observed sharpness on the step’s mini batch on the batch used for the step; Green: Batch Sharpness (Definition 1); Blue: full-batch $\lambda_{\max}$. Top row: MLP (2 hidden layers of width 512); middle: 5-layer CNN; bottom: ResNet-14; all trained on an 8k subset of CIFAR-10.
4.3 Batch Sharpness Governs EoSS
Following Cohen et al. [1] and the discussion in Section 3, we track how the training dynamics change when perturbing the hyper-parameters mid-training. Overall, we find that Batch Sharpness governs EoSS behavior—mirroring how $\lambda_{\max}$ operates in the full-batch EoS—while the full-batch $\lambda_{\max}$ lags behind or settles inconsistently, underlining the mini-batch nature of SGD stability.
Inducing catapults.
Increasing the step size $\eta$ or decreasing the batch size $b$ triggers a catapult spike in all the quantities in considerations and the training loss, before Batch Sharpness re-stabilizes near the updated threshold $2/\eta$, see Figures 7(a) and 7(c). This therefore pushes $\lambda_{\max}$ lower. Crucially, if stability was governed solely by $\lambda_{\max}$, this step-size or batch-size adjustment would have left the dynamics unchanged, as $\lambda_{\max}$ remains below both the original and revised $2/\eta$ thresholds.
Re-entering progressive sharpening.
Conversely, reducing $\eta$ raises the $2/\eta$ threshold. Conversely, increasing the batch size reduces the gap between Batch Sharpness and $\lambda_{\max}$. These changes prompt a new phase of progressive sharpening of Batch Sharpness, see Figures 7(b) and 7(d). Here, $\lambda_{\max}$ also rises, but ultimately stabilizes at a lower value than if the entire training had run with the smaller step size. Again, if stability was governed by $\lambda_{\max}$, this step-size adjustment would have had the same effect as starting from scratch with the new step size. Notably, Batch Sharpness initially decreases when reducing the learning rate—a counterintuitive result potentially linked to shifting alignments between gradients and Hessian eigenvectors, warranting further investigation.
@@ 4.3 Batch Sharpness Governs EoSS
Figure 7: Effects of changing step size or batch size in EoSS. (a) Increasing the learning rate causes a catapult spike before Batch Sharpness re-settles at the new $2/\eta$. (b) Decreasing the learning rate prompts renewed progressive sharpening. (c) Increasing the batch size lowers Batch Sharpness and re-starts progressive sharpening. (d) Decreasing the batch size increases Batch Sharpness and causes a catapult. The experiments are conducted on a 32k subset of CIFAR-10 to ensure sufficient complexity remains in the dataset, which is necessary for observing renewed progressive sharpening, consistent with observations by Cohen et al. [1].
5 On Stability
The previous section empirically demonstrated that mini-batch SGD generally settles into the EoSS regime, where Batch Sharpness hovers around $2/\eta$. But why does this threshold matter, and what does it mean for the algorithm’s stability? In classical (full-batch) gradient descent, the condition $\eta<2/\lambda_{\max}$ guarantees local stability by preventing divergence along the direction of the largest eigenvalue of a fixed Hessian. Conversely, the stability of mini-batch SGD is complicated by the variability of Hessians across iterations, making traditional stability criteria insufficient, see Appendix L.
In this section, we investigate what the notion of stability for SGD on the landscapes of neural networks is and clarify how mini-batch SGD’s “edge” behavior emerges from the evolving distribution of mini-batch Hessians aligned with their own gradients—rather than from any single, fixed Hessian.
5.1 Mathematizing Instability
For our analysis, the relevant definition of Type-2 stability must capture the scenario in which perturbations of the step size trigger significant shifts to different regions of the loss landscape—rather than merely resulting in marginal re-stabilization at slightly higher loss values. This perspective naturally leads to a curvature-based criterion that distinguishes genuinely disruptive oscillations from benign noise-driven fluctuations. Specifically, when mini-batch loss surfaces are locally well-approximated by quadratics along the directions of SGD steps, the nature of the oscillations becomes governed by these local quadratic approximations. Under such conditions, oscillations qualify as disruptive precisely if they would result in divergence when evaluated on the locally approximating quadratic landscape.
Definition 2 (Quadratic instability and Catapults).
Consider the quadratic approximation of all the data point landscapes $(\theta-x_{i})^{\top}\mathcal{H}_{i}(\theta-x_{i})$. We say that a step size is unstable if the trajectory exits all the compacts in which the quadratic approximation holds up to $\mathcal{O}(\eta)$ [^7] This means that either SGD seen as a linear dynamical systemis diverging or that the re-stabilization would happen be at a level which exits the largest region in which the quadratic approximation holds and so the dyanmics changes region.. We say the algorithm experienced a catapult when this event happened.
Under this definition, Type-1 oscillations are correctly classified as stable. Even if increasing the step size causes these oscillations to locally increase in size but re-stabilize at the new level, the divergence remains bounded, remaining within a compact region, see for instance Figures 3 and 4.
5.2 Stability and Batch Sharpness
The next theorem implies that SGD is quadratically unstable in two cases, when $\lambda_{\max}>2/\eta$, but also more surprisingly when $\lambda_{\max}<2/\eta$ but Batch Sharpness is bigger than $2/\eta$. While in the first case one can show very easily that the parameters explode, in the second setting we can not and we instead show that the second moment of the mini-batch gradients $\mathbb{E}[\lVert\nabla L_{B}\rVert^{2}]$ increases exponentially.
Theorem 1.
Let $\eta\leq 2/\lambda_{\max}$. There exists a absolute constant $c>0$ such that if Batch Sharpness is strictly bigger than $2/\eta+c\eta$ then $\mathbb{E}[\lVert\nabla L_{B}\rVert^{2}]$ locally increases exponentially with the SGD step and the trajectory is quadratically unstable (Definition 2).
The proof is a little involved but essentially relies on Jensen and Cauchy-Schwarz inequalities, see Appendix C and Proposition 5.
Path dependency and implications on progressive sharpening.
understanding the inner working and the phenomenology of progressive sharpening is a key open ingredient to understand stability and location of convergence of optimization algorithms. In principle, instability could emerge while moving toward infinity in any possible direction on the space. Previous work dealt with sufficient and necessary conditions for the stability of SGD in general, [29, 32], i.e., when does SGD become unstable given any possible notion of progressive sharpening. We show that the trajectory goes to infinity in precise directions, the fact that Batch Sharpness is the quantity to look at implies that progressive sharpening is acting in a certain way on the datapoint Hessians. Precisely, that progressive sharpening takes the trajectory in that place of the boundary of the high dimensional open set of stability such that $GNI=BatchSharpness=2/\eta$. Empirically, indeed, the iterates become unstable precisely because of Batch Sharpness reaching $2/\eta$, not because of, e.g., $\lambda_{\max}$ increasing but its variance remaining stable (which would keep Batch Sharpness $<GNI-constant$ until the instability threshold. This, we think, shows one more time why it is key for a deep learning optimization theory to advance to focus on path-dependent properties of the correct paths. Importantly, in Appendix B we also establish exactly when $GNI\geq\textit{Batch Sharpness}$. This fact depends only on the kurtosis of the gradients (which rarely changes much), $\mathbb{E}[\lVert\mathcal{H}(L_{B})-\mathcal{H}\rVert^{2}_{2}]$, and the alignment of the steps with the full-batch Hessian. We conjecture that Batch Sharpness reaches $GNI$ at $2/\eta$ by means of increasing $\mathbb{E}[\lVert\mathcal{H}(L_{B})-\mathcal{H}\rVert^{2}_{2}]$ more than the alignment of gradients with the top eigenvectors the full-batch Hessian. This would imply that progressive sharpening increases the variance of the Hessian over the batches at least as fast as $\lambda_{\max}$. This conjecture would agree with the red cloud amplifying around the green line in Figure 6.
Stability on the mini-batch step.
It is also worth commenting on maybe the most unsurprising property of EoSS: that, up to higher orders in the step size, the average single mini batch step is stable on the mini-batch landscape if and only if Batch Sharpness is below $2/\eta$. Curiously, however, we were not able to deduce properties of the stability of the whole system solely from this observation, Theorem 1 was proved unrelatedly.
Lemma 1.
$\mathbb{E}_{\begin{subarray}{c}\mathcal{P}_{b}\end{subarray}}[L_{B}(x_{t+1})-L_{B}(x_{t})]\leq O(\eta^{2})$ if and only if $\eta\leq 2/\textit{Batch Sharpness}$. If the mini-batch landscapes satisfy the PL condition, also the mini-batch gradients do not blow up $\mathbb{E}_{\begin{subarray}{c}\mathcal{P}_{b}\end{subarray}}[\lVert\nabla L_{B}(x_{t+1})\rVert^{2}-\lVert\nabla L_{B}(x_{t})\rVert^{2})]\leq O(\eta^{2})$.
The importance of stating this result is that Lemma 1 together with Theorem 1 imply that: the curvature-driven oscillations on the average mini-batch landscape imply full-batch loss diverges; the oscillations on the full-batch landscape which are stable on the mini-batch one do not. When the dynamics on the mini-batches is stable (Type-1), the dynamics is stable and progressive sharpening still can happen (see Figure 4) until Batch Sharpness reaches $2/\eta$.
6 On the fate of ${\lambda_{\max}}$
In this section we examine how $\lambda_{\max}$ behaves once EoSS is reached and clarify its relationship to Batch Sharpness. A key aspect of the original EoS analysis is, indeed, that the controlling quantity—the largest eigenvalue of the full-batch Hessian $\lambda_{\max}$—has an immediate geometric interpretation. There exists an extensive literature about $\lambda_{\max}$ size and role in neural networks, and it is a main ingredient of any proof of convergence. The EoSS picture replaces $\lambda_{\max}$ with Batch Sharpness, a statistic whose connection to generalization and role in optimization theory is largely unexplored.
(a) MLP, $\eta=0.01$
(b) MLP, $\eta=0.004$
(c) CNN, $\eta=0.03$
Figure 8: Stabilisation level of $\boldsymbol{\lambda_{\max}}$ across learning rates and architectures. Top: final-epoch $\lambda_{\max}$ vs. batch size. Bottom: log–log plots of the gap  $\,2/\eta-\lambda_{\max}\,$ for the same runs. All experiments use CIFAR-10 8k; further details in Appendix H.
6.1 Empirical facts
Below the phenomena we extensively observe in vision classification tasks trained with MSE, ablating on batch sizes, learning rates, architectures, datasets. See Figure 6 for a good reference of what generally goes on.
  • Fact 1:  Progressive Sharpening. $\lambda_{\max}$ increases at most as long as Batch Sharpness increases.
  • Fact 2:  Phase Transition. Once Batch Sharpness plateaus at $2/\eta$, $\lambda_{\max}$ stops increasing. If it moves, it only decreases from this time on.
  • Fact 3:  Path-dependence. If changes to hyper parameters are made, Batch Sharpness changes abruptly or restart growing and $\lambda_{\max}$ also changes. Stabilization of both happen as Batch Sharpness reaches $2/\eta$. The trajectory of $\lambda_{\max}$ is inconsistent with the size of hyper parameters (see Figure 7). The level of $\lambda_{\max}$ is path-dependent: it inherits the history of progressive sharpening up to the moment EoSS is reached.
  • Fact 4:  Smaller batches $\Rightarrow$ flatter minima. Across every setting we tested, reducing the batch size monotonically decreases the plateau level of $\lambda_{\max}$. This aligns with the long-standing empirical observation that smaller batches locate flatter minima see, e.g., Keskar et al. [2], Jastrzębski et al. [33]).
  • Fact 5:  A critical batch size marks the SGD $\to$ GD crossover. Each curve in Figure 8 exhibits a bend at $b\approx b_{c}(\eta)$: for $b<b_{c}$ the plateau falls rapidly with $b$, while for $b\!>\!b_{c}$ it flattens and approaches the full-batch value. This $b_{c}$ corresponds to the regime in which the mini-batch landscapes approximate well enough the full-batch landscape, restoring GD-like dynamics (Appendix H).
  • Fact 6:  No universal power law. From static analysis, one would expect a scaling $2/\eta-\lambda_{\max}=O(b^{-\alpha})$ for some $\alpha$, see Appendix I. The log–log plots (bottom row of Figure 8) show no robust straight-line behaviour, ruling out such law for any possible exponent $-\alpha$.
6.2 Implications and Open Questions
The findings above lead to the following main conclusions.
  1. (C1)
    $\lambda_{\max}$ is not the stability limiter for mini-batch training. Batch Sharpness governs EoSS; $\lambda_{\max}$ follows. $\lambda_{\max}$ is capped from above by the value it reaches at the phase transition characterized by Batch Sharpness reaching $2/\eta$. This and Facts 1—3 above imply that: {mdframed} The stabilization of $\lambda_{\max}$ is a by-product of EoSS, not the quantity that governs it.
  2. (C2)
    A theory of $\lambda_{\max}$ has to account for the correct progressive sharpening. By fixing the model and changing batch size $b$, the gap between the maximal eigenvalue of $\mathbb{E}\big{[}\lambda_{\max}(\mathcal{H}(L_{B}))\big{]}$ and $\lambda_{\max}=\lambda_{\max}(\mathcal{H})=\lambda_{\max}\mathbb{E}[\mathcal{H}(L_{B})]$ scales as $1/b$, see a proof in Appendix I. Any theory that keeps the parameter vector fixed and only varies $b$, or anyways leads to a power law, misses the path-dependent descent that determines where training arrives and where $\lambda_{\max}$ stabilizes. Facts 3 and 6 thus imply that analysis of $\lambda_{\max}$ is insufficient if it does not account for (1) the precise and correct effect of progressive sharpening on the higher moments of the Hessian and (2) the correct alignment between mini-batch steps and Hessians.
Quantifying the plateau of $\lambda_{\max}$ is thus still an (important) open problem. A complete account will require a dynamical theory through the progressive-sharpening phase and beyond. Not just properties at its endpoint as for full-batch methods.
Remark: Why not $\lambda_{\max}\mathbb{E}[\mathcal{H}(L_{B})]$.
Figure 9: Ranging different batch sizes.
One might hope that the largest eigenvalue of each mini-batch Hessian could act as a stochastic proxy for curvature. The reason is that the step is generally not aligned with the eigenvector of the largest eigenvalue and thus $\lambda_{\max}\mathbb{E}[\mathcal{H}(L_{B})]$ stabilizes right above $2/\eta$, however, the level of stabilization slightly changes with $b$—see Appendix J—unlike for Batch Sharpness.
7 Implications: How Noise-Injected GD Differs from SGD
SGD vs. Noisy Gradient Descent.
Figure 10: SGD vs. Noisy GD. Only noise preserving the mini-batch structure of SGD leads to $\lambda_{\max}$ plateauing below $2/\eta$ (akin to EoSS and as observed by [2]). Noise injection fails to reproduce this behavior even with the same covariance SGD’s.
  • Anisotropic Sampling Noise: Gaussian reweighting on the samples [51], which is different from SGD but maintains the mini-batch structure.
  • Diagonal Noise: Gaussian noise restricted to the diagonal part of the SGD noise covariance [52].
  • Isotropic Noise: Gaussian noise with isotropic covariance [52].
As shown in Figure 10 and Appendix D, only noise which maintains the higher moments of the Hessian(s) leads to an EoSS-like regime with $\lambda_{\max}$ stabilizing well below $2/\eta$. More generic (e.g., diagonal or isotropic) noise fails to reproduce this behavior. These experiments suggest that stability thresholds differ fundamentally between mini-batch SGD (governed by Batch Sharpness) and noise-injected GD (governed by $\lambda_{\max}$). Unsurprisingly, in the case in which the noise affects only the gradients—not the Hessians—indeed, EoSS comes for $\lambda_{\max}=2/\eta$ as for GD [29, 32]. Even in the quadratic setting, the appearance of Type-1 oscillations and GNI are not affected by the structure and distribution of the Hessian on the mini-batches, see Appendix A. The stability threshold, however, is. It depends on the Hessian’s higher moments, see Theorem 1 or [29, 32].
Challenges for SDE Modeling.
Classical analyses of neural network optimization often assume a single, static landscape: (i) Online perspective, modeling each step’s gradient as a noisy unbiased estimator of the expected gradient, or (ii) Offline perspective, treating the dataset as fixed and SGD as noisy GD on the empirical loss. In both views, it is the full-batch Hessian that supposedly drives curvature. Our results instead highlight that each update sees a Hessian $\mathcal{H}(L_{\mathbf{B}})$ that generally differs significantly from $\mathcal{H}$, leading to Batch Sharpness stabilizing at $2/\eta$ when $\lambda_{\max}$ is smaller.